OpenStudy (anonymous):

A swimmer finishes a race in 200secs with an average speed of 3ft per second. Find the average speed of the swimmer if it takes 120secs to finish the race 4ft? 5ft? 6ft? 7ft?

OpenStudy (anonymous):

5ft/sec is the correct answer avg. speed=total distance travelled/total time taken to cover that distance therefore 3=d/200 which gives d=600ft. now we have d=600ft and t=120s to cover that distance therefore avg. speed in this case=600/120 which gives average speed=5ft/s